# The Klein-Gordon equation on asymptotically Minkowski spacetimes: causal propagators

### v1

@@ 2.5. Hamiltonian flow and radial sets
$j_{\mathrm{sc},m_{1}+m_{2}-1,r_{1}+r_{2}-1}\left(i[A,B]\right)$ $=\langle t,z)\rangle^{-r_{2}}\langle\tau,\zeta\rangle^{-m_{2}}\cdot{}^{\mathrm{sc}}\!H_{a}(b)\,.$
@@ 2.7. Radial points estimates
Let $s,\ell^{\prime}\in\mathbb{R}$, $\ell^{\prime}>-1/2$, $\ell^{\prime}\in[\boldsymbol{\ell}-1/2,\boldsymbol{\ell})$. Then there are are $B,G\in{}^{\mathrm{sc}}\Psi^{0,0}$ with $\mathcal{R}^{\prime}\subset\operatorname{Ell}(B)$ and $\operatorname{WF}^{\prime}(B)\subset\operatorname{Ell}(G)$, such that: if $Gu\in H_{\mathrm{sc}}^{s-1/2,\ell^{\prime}}$ and $GP_{0}u\in H_{\mathrm{sc}}^{s-1,\boldsymbol{\ell}+1}$, then $Bu\in H_{\mathrm{sc}}^{s,\boldsymbol{\ell}}$, and for all $M,N$ there is $C>0$ such that
@@ 2.8. Weight functions, Fredholm estimates, and propagators
Now we take a microlocal partition of unity $\operatorname{Id}=B_{1}+B_{2}+B_{3}+B_{4}$, $B_{i}\in{}^{\mathrm{sc}}\Psi^{0,0}(X)$ with $\operatorname{WF}^{\prime}(B_{i})\subset O_{i}$.
@@
(2.37) $\lVert B_{4}u\rVert_{s,\boldsymbol{\ell}_{+}}\leq C\left(\lVert P_{0}u\rVert_{s-2,\boldsymbol{\ell}_{+}}+\lVert u\rVert_{s-1/2\boldsymbol{\ell}^{\prime}}\right)\leq C\left(\lVert P_{0}u\rVert_{s-1,\boldsymbol{\ell}_{+}+1}+\lVert u\rVert_{s-1/2,\boldsymbol{\ell}^{\prime}}\right).$
@@
$\lVert u\rVert_{1-s,-1-\boldsymbol{\ell}_{+}}\leq C\left(\lVert P_{0}^{*}\rVert_{-s,-\boldsymbol{\ell}_{+}}+\lVert u\rVert_{-s-1/2,\boldsymbol{\ell}^{\prime}}\right),$
we use the same chain of estimates, but with the roles of $\mathcal{R}^{p}$ and $\mathcal{R}^{f}$ exchanged. In other words, we propagate regularity from $\mathcal{R}^{f}$ to $\mathcal{R}^{p}$ along the Hamiltonian flow in $\operatorname{Char}(P_{0})$. Here $\boldsymbol{\ell}^{\prime}$ must be chosen analogously, i.e., $\boldsymbol{\ell}^{\prime}<-1-\boldsymbol{\ell}_{+}$ must also be greater than the threshold $-1/2$ near $\mathcal{R}^{f}$. By the standard argument of iterating by $1/2$ differential orders we replace the $-1/2$ on the right by an arbitrary differential order $-N$ and deduce (2.33). Note that the formulation of the estimate in (2.33) with the lower order error term on the right hand side follows from bounding the $\boldsymbol{\ell}^{\prime}$ term by a small factor times that $\boldsymbol{\ell}$ term on the left; specifically, using that for any $M>0$ and $\epsilon>0$ there is $C(\epsilon)>0$ such that $x^{-\boldsymbol{\ell}^{\prime}}\leq C(\epsilon)x^{M}+\epsilon x^{\boldsymbol{\ell}}$, based off which we have, for any $N\in\mathbb{R}$, a $C>0$ such that
(2.38) $\lVert u\rVert_{H_{\mathrm{sc}}^{-N,\boldsymbol{\ell}^{\prime}}}\leq C\lVert u\rVert_{H_{\mathrm{sc}}^{-N,-M}}+\epsilon\lVert u\rVert_{H_{\mathrm{sc}}^{-N,-\boldsymbol{\ell}}},$
@@
With these estimates, we have nearly proved Theorem 2.1:.
@@ 3.2. $3\mathrm{sc}$-differential operators
$\operatorname{Diff}_{3\mathrm{sc}}^{m,l}=\langle t,z\rangle^{l}\operatorname{Diff}_{3\mathrm{sc}}^{m},$
@@ 3.3. $3\mathrm{sc}$-geometry
(3.12) $\pi\colon{}^{3\mathrm{sc}}T^{*}_{C}X\longrightarrow W^{\perp}\,,$
@@ 4.1. $3\mathrm{sc}$-pseudodifferential operators
(4.1) ${}^{3\mathrm{sc}}\!S^{m,r}(X;C)\ =\rho_{\operatorname{fib}}^{-r}\rho_{\infty}^{-m}C^{\infty}({}^{\mathrm{sc}}\overline{T}^{*}[X;C])$
@@ 4.2. The principal symbol and the indicial operator
$\hat{A}_{\operatorname{ff},\ell}\coloneqq\hat{N}_{\operatorname{ff},\ell}(A)\,.$
@@
(4.9) $a_{\operatorname{ff}}\coloneqq(x^{r}a)\rvert_{{}^{3\mathrm{sc}}\overline{T}_{\operatorname{ff}}X},$
@@
(4.12) $j_{3\mathrm{sc},m,r}:{}^{3\mathrm{sc}}\Psi^{m,r}$ $\longrightarrow C^{\infty}({}^{3\mathrm{sc}}\overline{T}_{\operatorname{mf}}X)\times C^{\infty}({}^{3\mathrm{sc}}\!S^{*}[X;C])\times C^{\infty}(\mathbb{R}_{\tau};{}^{\mathrm{sc}}\Psi^{m,0}(\operatorname{ff}))$
@@
The kernel of the mapping of the principal symbol mapping (4.12) is exactly ${}^{3\mathrm{sc}}\Psi^{m-1,r-1}$, while the image is the set of those $(q_{1},q_{2},\{Q_{\tau}\})$ such that, with $q_{0}$ the left reduction of $Q_{\tau}$, we have $\langle\tau,\zeta\rangle^{-m}q_{0}\in C^{\infty}(\operatorname{ff}\times\overline{\mathbb{R}^{n+1}_{\tau,\zeta}})$ and the matching condition (4.15) hold. Moreover, the mapping $j_{3\mathrm{sc}}$ is multiplicative (but not commutative): for $A\in{}^{3\mathrm{sc}}\Psi^{m_{1},r_{1}},B\in{}^{3\mathrm{sc}}\Psi^{m_{2},r_{2}}$
@@ 4.5. Wavefront sets and elliptic sets
(4.27) ${}^{\mathrm{sc}}\dot{\overline{T}}X=\left({}^{\mathrm{sc}}\overline{T}^{*}X\setminus{}^{\mathrm{sc}}\overline{T}^{*}_{C}X\right)\cup\overline{W^{\perp}},$
@@
${}^{\mathrm{sc}}\dot{T}X=\left({}^{\mathrm{sc}}T^{*}X\setminus{}^{\mathrm{sc}}T^{*}_{C}X\right)\cup W^{\perp},$
@@ 5.1. Functional calculus
$(\widehat{G_{\psi}})_{\operatorname{ff}}-(\widehat{G_{\psi,0}})_{\operatorname{ff}}\in\Psi_{\mathrm{scl},\mathrm{sc}}^{-\infty,-1,-1}\,.$
@@
$(\widehat{G_{\psi}})_{\operatorname{ff}}(\tau)=\psi\left((\tau^{2}+H_{V_{0}}+E)^{-1}(\tau^{2}-H_{V_{0}})\right)\,.$
@@
$j_{\mathrm{sc}}\left((\widehat{G_{\psi}})_{\operatorname{ff}}(\tau)\right)=j_{\mathrm{sc}}\left((\widehat{G_{\psi,0}})_{\operatorname{ff}}(\tau)\right)$
@@
$\sigma_{\mathrm{scl},h=1/\tau}\left((\widehat{G_{\psi}})_{\operatorname{ff}}\right)=\psi\left((1+h^{2}\Delta)^{-1}(1-h^{2}\Delta)\right)\,,$
@@
5.2. Further elliptic regularity near $C$
If $P_{V}u\in H_{\mathrm{sc}}^{s,\ell}$, then away from $C$, then on ${}^{3\mathrm{sc}}\!\operatorname{Ell}(P_{V})$ we have the estimates in Section 4.6. At $C$, frequency localization becomes global in the interaction variables, and, as discussed in above, we use $G_{\psi}$ to localize near the characteristic set. Thus, for $\psi\in C^{\infty}_{c}$ with $\psi(s)=1$ for $|s|<\delta$, $\delta>0$, the operator
$\operatorname{Id}-G_{\psi}$
is morally speaking a localizer to an elliptic region over $\mathrm{NP}$ (or $\mathrm{SP}$). Indeed, for $P_{V}$ with $V$ non-static, if $Q$ is a localizer to a neighborhood of ${}^{\mathrm{sc}}\overline{T}^{*}_{\mathrm{NP}}X$ and $G_{\psi}$ is defined using $V_{+}$, then $Q(\operatorname{Id}-G_{\psi})$ is a reasonable generalization of localization to the elliptic set from the $\mathrm{sc}$ setting.
Thus we expect to have an elliptic regularity statement using $\operatorname{Id}-G_{\psi}$, and such a statement will be crucial below as all the propagation estimates are proven first using commutators which have factors of $G_{\psi}$ to localize to the characteristic set. Specifically, if $P_{V}u\in H_{\mathrm{sc}}^{s,\ell}$ then we should have $(\operatorname{Id}-G_{\psi})u\in H_{\mathrm{sc}}^{s+2,\ell}$ as well, at least locally near $poles$. That is what we prove in this section.
We do so with a parametrix construction Note first that for static $V=V_{0}$, we have
(5.12) $E_{\psi}P_{V_{0}}=\operatorname{Id}-G_{\psi}$
on the nose, since, if
(5.13) $E_{\psi}=F_{\psi}\circ(D_{t}^{2}+H_{V_{0}}+E)^{-1}$
we can set
$F_{\psi}=f_{\psi}((D_{t}^{2}+H_{V_{0}}+E)^{-1}P_{V_{0}})\quad\mbox{ with }\quad f_{\psi}(s)=\frac{1-\psi(s)}{s}.$

### v2

@@ 2.5. Hamiltonian flow and radial sets
$j_{\mathrm{sc},m_{1}+m_{2}-1,r_{1}+r_{2}-1}\left(i[A,B]\right)$ $=\langle t,z\rangle^{-r_{2}}\langle\tau,\zeta\rangle^{-m_{2}}\cdot{}^{\mathrm{sc}}\!H_{a}(b)\,.$
@@ 2.7. Radial points estimates
Let $s,\ell^{\prime}\in\mathbb{R}$, $\ell^{\prime}>-1/2$, $\ell^{\prime}\in[\boldsymbol{\ell}-1/2,\boldsymbol{\ell})$. Then there are $B,G\in{}^{\mathrm{sc}}\Psi^{0,0}$ with $\mathcal{R}^{\prime}\subset\operatorname{Ell}(B)$ and $\operatorname{WF}^{\prime}(B)\subset\operatorname{Ell}(G)$, such that: if $Gu\in H_{\mathrm{sc}}^{s-1/2,\ell^{\prime}}$ and $GP_{0}u\in H_{\mathrm{sc}}^{s-1,\boldsymbol{\ell}+1}$, then $Bu\in H_{\mathrm{sc}}^{s,\boldsymbol{\ell}}$, and for all $M,N$ there is $C>0$ such that
@@ 2.8. Weight functions, Fredholm estimates, and propagators
Now we take a microlocal partition of unity $\operatorname{Id}=B_{1}+B_{2}+B_{3}+B_{4}$, $B_{i}\in{}^{\mathrm{sc}}\Psi^{0,0}(X)$ with
(2.34) $\operatorname{WF}^{\prime}(B_{i})\subset O_{i}.$
@@
(2.38) $\lVert B_{4}u\rVert_{s,\boldsymbol{\ell}_{+}}\leq C\left(\lVert P_{0}u\rVert_{s-2,\boldsymbol{\ell}_{+}}+\lVert u\rVert_{s-1/2,\boldsymbol{\ell}^{\prime}}\right)\leq C\left(\lVert P_{0}u\rVert_{s-1,\boldsymbol{\ell}_{+}+1}+\lVert u\rVert_{s-1/2,\boldsymbol{\ell}^{\prime}}\right).$
@@
$\lVert u\rVert_{1-s,-1-\boldsymbol{\ell}_{+}}\leq C\left(\lVert P_{0}^{*}u\rVert_{-s,-\boldsymbol{\ell}_{+}}+\lVert u\rVert_{-s-1/2,\boldsymbol{\ell}^{\prime}}\right),$
we use the same chain of estimates, but with the roles of $\mathcal{R}^{p}$ and $\mathcal{R}^{f}$ exchanged. In other words, we propagate regularity from $\mathcal{R}^{f}$ to $\mathcal{R}^{p}$ along the Hamiltonian flow in $\operatorname{Char}(P_{0})$. Here $\boldsymbol{\ell}^{\prime}$ must be chosen analogously, i.e., $\boldsymbol{\ell}^{\prime}<-1-\boldsymbol{\ell}_{+}$ must also be greater than the threshold $-1/2$ near $\mathcal{R}^{f}$.
By the standard argument of iterating by $1/2$ differential orders we replace the $-1/2$ on the right by an arbitrary differential order $-N$ and deduce (2.33). Note that the formulation of the estimate in (2.33) with the lower order error term on the right hand side follows from bounding the $\boldsymbol{\ell}^{\prime}$ term by a small factor times that $\boldsymbol{\ell}$ term on the left; specifically, using that for any $M>0$ and $\epsilon>0$ there is $C(\epsilon)>0$ such that $x^{-\boldsymbol{\ell}^{\prime}}\leq C(\epsilon)x^{M}+\epsilon x^{-\boldsymbol{\ell}}$, based off which we have, for any $N\in\mathbb{R}$, a $C>0$ such that
(2.39) $\lVert u\rVert_{-N,\boldsymbol{\ell}^{\prime}}\leq C\lVert u\rVert_{-N,-M}+\epsilon\lVert u\rVert_{-N,\boldsymbol{\ell}}\,,$
@@
With these estimates, we have nearly proved Theorem 2.1:
@@ 3.2. $3\mathrm{sc}$-differential operators
$\operatorname{Diff}_{3\mathrm{sc}}^{m,r}=\langle t,z\rangle^{r}\operatorname{Diff}_{3\mathrm{sc}}^{m},$
@@ 3.3. $3\mathrm{sc}$-geometry
(3.12) $\pi\colon{}^{\mathrm{sc}}T^{*}_{C}X\longrightarrow W^{\perp}\,,$
@@
We can now extend the defintion of the radial set to the $3\mathrm{sc}$-setting for $P_{0}$. It is given by
(3.14) ${}^{3\mathrm{sc}}\mathcal{R}\coloneqq\mathcal{R}\cap\left({}^{\mathrm{sc}}S^{*}_{X\setminus C}X\cup{}^{\mathrm{sc}}T^{*}_{\partial X\setminus C}X\right)\cup\left(C\times\{\pm m\}\right)$
and
${}^{3\mathrm{sc}}\mathcal{R}_{\operatorname{src}}$ $\coloneqq\mathcal{R}_{\operatorname{src}}\cap\left({}^{\mathrm{sc}}S^{*}_{X\setminus C}X\cup{}^{\mathrm{sc}}T^{*}_{\partial X\setminus C}X\right)\cup\mathcal{R}_{\operatorname{ff},\operatorname{src}}\,,$
${}^{3\mathrm{sc}}\mathcal{R}_{\operatorname{snk}}$ $\coloneqq\mathcal{R}_{\operatorname{snk}}\cap\left({}^{\mathrm{sc}}S^{*}_{X\setminus C}X\cup{}^{\mathrm{sc}}T^{*}_{\partial X\setminus C}X\right)\cup\mathcal{R}_{\operatorname{ff},\operatorname{snk}}\,.$
We have that ${}^{3\mathrm{sc}}\mathcal{R}={}^{3\mathrm{sc}}\mathcal{R}_{\operatorname{src}}\sqcup{}^{3\mathrm{sc}}\mathcal{R}_{\operatorname{snk}}$ and ${}^{3\mathrm{sc}}\mathcal{R}_{\bullet}\subset\gamma_{3\mathrm{sc}}(\mathcal{R}_{\bullet})$ for $\bullet\in\{\operatorname{src},\operatorname{snk}\}$.
@@ 4.1. $3\mathrm{sc}$-pseudodifferential operators
(4.1) ${}^{3\mathrm{sc}}\!S^{m,r}(X;C)\ =\rho_{\operatorname{fib}}^{-m}\rho_{\infty}^{-r}C^{\infty}({}^{\mathrm{sc}}\overline{T}^{*}[X;C])$
@@ 4.2. The principal symbol and the indicial operator
$\hat{A}_{\operatorname{ff},r}\coloneqq\hat{N}_{\operatorname{ff},r}(A)\,.$
@@
(4.9) $a_{\operatorname{ff}}\coloneqq(x^{r}a)\rvert_{{}^{3\mathrm{sc}}\overline{T}_{\operatorname{ff}}[X;C]},$
@@
(4.12) $j_{3\mathrm{sc},m,r}:{}^{3\mathrm{sc}}\Psi^{m,r}$ $\longrightarrow C^{\infty}({}^{3\mathrm{sc}}\overline{T}_{\operatorname{mf}}^{*}[X;C])\times C^{\infty}({}^{3\mathrm{sc}}\!S^{*}[X;C])\times C^{\infty}(\mathbb{R}_{\tau};{}^{\mathrm{sc}}\Psi^{m,0}(\operatorname{ff}))$
@@
The kernel of the principal symbol mapping (4.12) is exactly ${}^{3\mathrm{sc}}\Psi^{m-1,r-1}$, while the image is the set of those $(q_{1},q_{2},\{Q_{\tau}\})$ such that, with $q_{0}$ the left reduction of $Q_{\tau}$, we have $\langle\tau,\zeta\rangle^{-m}q_{0}\in C^{\infty}(\operatorname{ff}\times\overline{\mathbb{R}^{n+1}_{\tau,\zeta}})$ and the matching condition (4.15) hold. Moreover, the mapping $j_{3\mathrm{sc}}$ is multiplicative (but not commutative): for $A\in{}^{3\mathrm{sc}}\Psi^{m_{1},r_{1}},B\in{}^{3\mathrm{sc}}\Psi^{m_{2},r_{2}}$
@@ 4.5. Wavefront sets and elliptic sets
(4.27) ${}^{\mathrm{sc}}\dot{\overline{T}}^{*}X=\left({}^{\mathrm{sc}}\overline{T}^{*}X\setminus{}^{\mathrm{sc}}\overline{T}^{*}_{C}X\right)\cup\overline{W^{\perp}},$
@@
${}^{\mathrm{sc}}\dot{T}^{*}X=\left({}^{\mathrm{sc}}T^{*}X\setminus{}^{\mathrm{sc}}T^{*}_{C}X\right)\cup W^{\perp},$
@@ 5.1. Functional calculus
Moreover, we have an elliptic estimate for $BG_{\varphi}u$ by $QG_{\psi}u$ given that $Q$ is elliptic on the wavefront set of $BG_{\varphi}$ and the support of $\varphi$ is contained in the set $\{\psi=1\}$.
Proposition 5.8.
Let $B,Q\in{}^{3\mathrm{sc}}\Psi^{m,r}$ and $\varphi,\psi\in C_{c}^{\infty}(\mathbb{R})$ with $\varphi\psi=\varphi$ and $\varphi(0)=1$. If ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(BG_{\varphi})\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(Q)$, then for any $N,M\in\mathbb{N}$ there exists $C>0$ such that for $u\in H_{\mathrm{sc}}^{-N,-M}$,
$\lVert BG_{\varphi}u\rVert\leq C\left(\lVert QG_{\psi}u\rVert+\lVert u\rVert_{-N,-M}\right)\,.$
Proof.
Using $3\mathrm{sc}$-elliptic regularity for $BG_{\varphi}$, then using $G_{\varphi}G_{\psi}=G_{\varphi}$,
$\lVert BG_{\varphi}u\rVert=\lVert(BG_{\varphi})G_{\psi}u\rVert\leq C\left(\lVert QG_{\psi}u\rVert+\lVert u\rVert_{-N,-M}\right),$
as claimed. ∎
@@
$\hat{N}_{\operatorname{ff}}(G_{\psi})-\hat{N}_{\operatorname{ff}}(G_{\psi,0})\in\Psi_{\mathrm{scl},\mathrm{sc}}^{-\infty,-1,-1}\,.$
@@
$\hat{N}_{\operatorname{ff}}(G_{\psi})(\tau)=\psi\left((\tau^{2}+H_{V_{0}}+E)^{-1}(\tau^{2}-H_{V_{0}})\right)\,.$
@@
$j_{\mathrm{sc}}\left(\hat{N}_{\operatorname{ff}}(G_{\psi})(\tau)\right)=j_{\mathrm{sc}}\left(\hat{N}_{\operatorname{ff}}(G_{\psi,0})(\tau)\right)$
@@
$\sigma_{\mathrm{scl},h=1/\tau}\left(\hat{N}_{\operatorname{ff}}(G_{\psi})\right)=\psi\left((1+h^{2}\Delta)^{-1}(1-h^{2}\Delta)\right)\,,$
@@
Definition 5.13.
Let $V\in\rho_{\operatorname{mf}}{}^{3\mathrm{sc}}\Psi^{1,0}$ and $r\in(0,\infty)$. We say that $V$ is asymptotically static of order $r$ at $C$, if there exist $V_{\pm}\in S_{\mathrm{cl}}^{-1}(\mathbb{R}^{n}_{z})$ and $\chi_{\pm}\in C_{c}^{\infty}(X)$ such that
  1. (1)
    $\chi_{+}(\mathrm{NP})=1$ and $\chi_{-}(\mathrm{SP})=1$,
  2. (2)
    $\chi_{\pm}\cdot(V-V_{\pm})\in{}^{3\mathrm{sc}}\Psi^{1,-r}$.
We note that if $V$ is asymptotically static of order $r>0$, then the static parts $V_{\pm}$ are uniquely determined.
Because the operators $G_{\varphi}$ act as localizers to the characteristic set of $P_{V}$ over the north pole, $Bu$ should morally be controlled by $BG_{\varphi}u$ and $P_{V}u$. We now prove an estimate in this direction when $B\in{}^{3\mathrm{sc}}\Psi^{s,\ell}$ is localized near $\mathrm{NP}$. (An equivalent estimate holds near $\mathrm{SP}$.) We bound $Bu$ in terms of $BG_{\psi}u$ and $GP_{V}u$ if $G$ is elliptic on the wavefront set of $B$. This estimate is used extensively below to convert control of $BG_{\varphi}u$ (which is convenient to obtain via our propagation estimates) into control of $Bu$ when $B$ is elliptic (the latter control is harder to obtain because any elliptic set over the front face necessitates that the entire fiber over the front face is in the operatorial wavefront set):
Proposition 5.14.
Let $V\in\rho_{\operatorname{mf}}{}^{3\mathrm{sc}}\Psi^{1,0}$ be asymptotically static of order $r$. Let $\varphi\in C_{c}^{\infty}(\mathbb{R})$ with $\varphi|_{(-\varepsilon,\varepsilon)}\equiv 1$ for some $\varepsilon>0$, $B,G^{\prime},B^{\prime}\in{}^{3\mathrm{sc}}\Psi^{0,0}$ such that $B$ is microsupported near $\mathrm{NP}$ and
${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(B)\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(G^{\prime})\cap{}^{3\mathrm{sc}}\!\operatorname{Ell}(B^{\prime})\,.$
For all $M,N\in\mathbb{N}$ and $s,\ell\in\mathbb{R}$ there exists $C>0$ such that for all $u\in H_{\mathrm{sc}}^{-N,-M}$,
$\lVert Bu\rVert_{s,\ell}\leq C\left(\lVert BG_{\varphi}u\rVert_{s,\ell}+\lVert G^{\prime}P_{V}u\rVert_{s-2,\ell}+\lVert B^{\prime}u\rVert_{s-1,\ell-r}+\lVert u\rVert_{-N,-M}\right)\,,$
provided that the right hand side is finite.
The main idea behind this proposition, which is proved below, is that $I-G_{\varphi}$ should act as a localizer to the elliptic set of $P_{V}$. This requires explanation, as it is in fact not such a localizer, i.e. ${}^{3\mathrm{sc}}\!\operatorname{WF}^{\prime}(I-G_{\varphi})\not\subset{}^{3\mathrm{sc}}\!\operatorname{Ell}(P_{V})$. However, near $\mathrm{NP}$, the leading order part of the operator, $P_{V_{+}}$, admits an explicit parametrix with error $G_{\varphi}$. That is, there exists $E_{\varphi}\in{}^{3\mathrm{sc}}\Psi^{-2,0}$ solving
(5.13) $E_{\varphi}P_{V_{+}}=\operatorname{Id}-G_{\varphi}.$
Indeed, this is satisfied by the operator
$E_{\varphi}\coloneqq F_{\varphi}\circ(D_{t}^{2}+H_{V_{+}}+E)^{-1}$
where
$F_{\varphi}=f_{\varphi}((D_{t}^{2}+H_{V_{+}}+E)^{-1}P_{V_{+}})\quad\mbox{ with}\quad f_{\varphi}(s)=\frac{1-\varphi(s)}{s}.$